OpenStudy (anonymous):

A small sphere with mass 5.00.×10−7kg and charge +3.00μC is released from rest a distance of 0.600 m above a large horizontal insulating sheet of charge that has uniform surface charge density σ=+8.00pC/m2. a)Using energy methods, calculate the speed of the sphere when it is 0.250 m above the sheet of charge? I just need help understanding what I have to do to get the speed. I tried finding the difference in potential to calculate the Work done and then find the final velocity from that. I tried using conservation of energy using Electric Potential energy, gravitational PE and KE.

OpenStudy (irishboy123):

devil might be in detail. what are you using for the electrical field of the sheet?

OpenStudy (anonymous):

use Gauss' Law to obtain E above the sheet; is E up, or is E down? Is the Electric Force up, or down?

OpenStudy (michele_laino):

the situation described in your problem is: 

OpenStudy (michele_laino):

of course, we can neglect the weight force of our point-like charge. Furthermore, we have the subsequent equations: \[\Large \begin{gathered} v = \frac{{qE}}{m}t \hfill \\ \hfill \\ z = {z_0} + \frac{1}{2}\frac{{qE}}{m}{t^2} \hfill \\ \end{gathered} \] where: \[\Large \begin{gathered} q = 3 \times {10^{ - 6}}coulombs \hfill \\ \\ E = \frac{\sigma }{{{\varepsilon _0}}} \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

here, you have to solve the second equation for t, namley: \[\Large t = \sqrt {\frac{{2m}}{{qE}}\left( {z - {z_0}} \right)} \] where z=0.250 meters then you have to substitute the resultant value, into the first equation: \[\Large v = \frac{{qE}}{m}t\]

OpenStudy (michele_laino):